Grade 12Algebra

Construct 3 equations starting with x = – 2.........

Pawan Prajapati
5 Years agoGrade 12

1 Answer

Harshit Singh
5 Years ago
Dear Student

First equation is,
Multiply both side by 5
= 5x = -10 ...[equation 1]
Second equation is,
Subtracting 3 from both side,
= 5x–3 = - 10–3
= 5x - 3 = - 13 ... [equation 2]
Third equation is,
Dividing both sides by 2
= (5x/2)–(3/2) = (-13/2) ... [equation 3]

Thanks